OpenStudy (anonymous):

Triangle ABC has vertices A(0,0), B(2,2) and (5,-1) Find the coordinates of L, the midpoint of AC and M the midpoint of BC. Verify LM is parallel to AB and LM=1/2 AB.

OpenStudy (amity):

The mid point of (x1,y1),(x2,y2),,,,,is x1+x2//2,andy1+y2/2

OpenStudy (amity):

The mid point of (x1,y1),(x2,y2),,,,,is x1+x2//2,andy1+y2/2

OpenStudy (amity):

L= (0+5,/2,0-1/2)=(5/2,-1/2)

OpenStudy (anonymous):

thank you so much!!